FEL Monthly Brief
October 2001

We began the last operational period for the IR Demo this month.   The FEL was operated for carbon nanotube
production, chemical dynamics, measurements of the optical pulse length, wiggler efficiency measurements, and
Compton x-ray production.  The first evidence for FEL excitation of a molecule in a molecular beam was
obtained by scientists from the UVA Chemistry Department.  Future experiments will provide valuable insight
into chemical dynamics.

Results from the optical output characterization measurements show that the optical pulse length can be
shortened to as low as 180 fs (FWHM).  We continued our measurements of high efficiency (> 1%) wiggler
operation and further  characterized the optical beam quality by making beam profile measurements (M2) as a
function of power.  We continued our characterization measurements on the wiggler for the IR Upgrade. We
are pleased with wiggler characteristics measured thus far.

The drawings and manufacturing specifications for the largest magnets in the Upgrade, the 180° recirculation
dipoles (DY),were signed and the package was sent to Procurement.

The 3rd 100 kW was received from the vendor this week, which completes this contract.  The 1st 100 kW klystron
and solenoid were installed on their cart and installation of the LCW was started.  Vacuum of all three (3) klystrons
proved to be good.

Members of the FEL project management team spent time with key managers and group leaders in the
Accelerator Division in a retreat this month to examine our present and future business lines.  Much of our time
was spent analyzing our capability to complete the 10 kW and 1 kW upgrade projects, the organization needed to
minimize conflicts and maximize benefits of our related projects in the CEBAF Upgrade and the SNS
accelerator, and the opportunities afforded by the potential 100 kW upgrade project and other future light source

A memo was distributed to the review team and invited guests for the next semiannual review for the FEL
Upgrade project on Nov. 8-9, soliciting particular agenda topics to supplement the project status reports.  We
had an extended teleconference with the ONR Contract monitor  to review the close-out of the Phase 1
contract on the FEL Upgrade project which will be reviewed in detail during the project semiannual review on
Nov. 8-9th.  The agenda for the semiannual review of the FEL Upgrade project was reviewed with the ONR
contract monitor and distributed to the review team and prospective attendees.   Members of FEL team
continued their preparations for presenting summaries of the technical, cost and schedule status of each
Upgrade subsystem at the review.

A budget proposal for FY03-04 was prepared for the Commonwealth of Virginia for support of FEL operations
and an addition to the FEL Building to house the Helios synchrotron with additional laboratory and support
space.  These proposals are due at the Commonwealth Dept. of Planning and Budget on Oct. 14th.

We were alerted that the High Energy Laser panel of the Defense Science Board has released their report this
week which calls for an increase in DOD spending for laser R&D from $100M/yr to $150M/yr-including work
on free electron lasers.

The JLab and Aerospace procurement organizations successfully completed negotiations for our major
subcontract under the first phase of the UV FEL project. The package was forwarded to DOE in November
for approval.

Bill Nay, who oversees security issues for DOE's Office of Science, visited the Lab on Monday (Oct. 15th) to
discuss potential security/export control implications of the lab's activities in the design and operation of high
average power free electron lasers.

We completed the project monthly reports (technical and cost-performance) and distributed them to the ONR
and DOE contract  monitors.

We had two visitors in October from Dahlgren who are working on cryogenic systems for the “All Electric
Ship”, Dr. E. Michael Golda, Head of the Electrical Systems Department and Thomas Fikse, Lead Engineer.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September.  The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.  Costs accrued to Phase 1 during this period are bookings for previously committed funds as equipment is received.  When the open commitments are fully costed on or before December 30, 2001, we predict that the remaining Phase 1 cost and schedule variances will be approximately equal to -$1,300k.  Accordingly, the final Phase 1 cost performance report is planned for January, 2002. We expect to recover remaining Phase 1 cost and schedule variances by the completion of the Phase 2 barring unforeseen difficulties.

The Phase 2 project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.
Work for Phase 2 proceeded to plan during this reporting period.  Actual charges of $297k were accrued
for October for total accrued costs of  $934k to date for Phase 2.  The bulk of this effort involves long term
procurements of components designed during Phase 1.  No significant earned value is planned for the first
quarter of FY02.  We expect to meet the planned cost at completion of Phase 2 effort.  At present the cost
variance for Phase 2 is -$919k and the schedule variance is -$52k.  The program is presently on track to begin
commissioning efforts (presently un-funded) in September, 2002.

The UV FEL project was approved for $2,735k for a performance period from September 2001 to September
30, 2002.  Efforts were initiated last month with the official project start toward the project deliverables:  design
of the optical cavity system, design of the UV FEL electron beam and optical transport systems, and production
of harmonically generated UV light.  Significant earned value credit should be available by the end of the first
quarter of FY02 for all of these activities.

WBS 3 (Beam Physics):
IR Upgrade Driver Design Revision 1.1.2 was completed.  Work started on UV Upgrade Driver Design
Revision 1.1.2.  An entertaining feature of energy recovery during multipass operation of the Demo driver was
documented as JLAB-TN-01-048, and is available on line at:  http://www.jlab.org/~douglas/FEL/technote/JLABTN01048.pdf

Not only is it possible to simultaneously manage momentum spread on the coasting and energy recovery passes
one can also (for a "magic phase") image the injected beam at the dump.

DIMAD was ported to WinNT and is being tested.  BBU threshold simulations are being set up for the Revision
1.1.2 design. Documentation of the Revision 1.1.2 design is nearly complete.

BBU threshold simulations are being set up for the Revision 1.1.2 design.

GX and GQ magnet design analyses were reviewed; though the behaviors were largely acceptable, an inboard
field roll-off of order  0.03% observed in both magnets is of some concern and is being examined more

Correction and diagnostic configurations for the linac-to-endloop match were considered; magnetic
measurements were discussed.

WBS 4 (Injector):
Leak checked the ceramic stack for the upgrade gun and found two leaks in one kovar to ceramic braze. The
braze joint was cut off the stack and preparations for welding a new one on are being done. The shop received
the cathode support tube for machining.  The port tubes for the gun chamber were machined and

The drawings for the modified cathode shield design were submitted to the shop for manufacturing. Drawings
for the shield operating linkage are completed and under review.

A test section of the wire for feeding power to the cesium channels inside the ball cathode was made and works
very well.  The drawings are being fine tuned prior to sign off.  Concepts of a new resistor design allowing for
the removal of the elbow from the high voltage transmission line (would increase reliability) are being looked at.
Plans are starting to move the HVPS tank to be able to eliminate the elbow and the transmission line hangers.

A stainless steel reference electrode was treated in the newly re-commissioned ion implantation system at W&M.

WBS 5 (SRF):
Received second space frame and vacuum vessel from the vendor.  These include fixes and modifications as
identified in the first article.  Inspection and acceptance is in process.

WBS 6 (RF):
Zone 3 - Operating the zone into shorts.  Calibration curves have been run on all 8 klystron/ cavity pairs. The
final conformal cables need to be installed.

Zone 4 - All of the RF Control Modules have been calibrated and installed in the zone.  New CAMAC Stepper
Cards have been ordered to replace the missing ones.  They will be here in November.

Quarter Klystrons- Equipment and hardware is being collected in order to mount the first solenoid and 100 kW
klystron.  A lifting plate is being designed for the klystron.  The vendor, NWL, was contacted about making the
10 H smaller.  A "Hold" was placed on its manufacture while its design is being checked. The 3rd 100 kW
klystron has been tested, approved, shipped and received.  This completes the CPI contract. The klystron lift
plate drawing is signed, built, and tested.  The LCW manifolds were received for the circulators, RF loads, and
HOM filters.  Weekly meetings have started on the upgrade of the HVPS and klystron carts.

Quarter HVPS - The design is at the final stage of finding vendors for practical inductors and capacitors.  The
rough layout of the HVPS components is in progress.

Injector RF - Ion pump controllers were ordered this week for the 100 kW klystrons.  The lifting plate is
designed and parts are on order.  The designed will be approved before the part is built.

Injector HVPS - The design of the HVPS is nearly complete with the specification of the HV inductor and
capacitors.  Quotes are coming in for the purchase of the capacitors.  The layout of the HVPS major
components is in progress.

WBS 8 (Instrumentation):
Data Bases for the components of the Demo machine and the Upgrade machine are being developed for
tracking  removal and storage of components as well as their subsequent re-installation. Bar Codes are being
developed for component tracking during the Demo rip-out and subsequent installation of new equipment.
Material and equipment is being inventoried and staged for use and will be included in the data base as the
system is developed.

Firewire cables have been run to User Lab 1 and User Lab 3 to support new software using this technology. 5
new IR cameras have been modified and tested, 1 had a "twitch". One of those has been installed at Can 2 for
beam diagnostics at that location.  The new Areotech translation stages are being modified to conform with the
existing stepper motor interface chassis.

The new digital IR cameras are being commissioned. These are used as optical beam position monitors for the
FEL transport line.  They have both analog and digital outputs, the power and analog connections have been
broken out for immediate use. Digital control cables are on order and will arrive next week. The 30Hz sync and
distribution system is operational and ready for testing when the control cable arrive.

The four channel remote relay box is fully operational and documented. It allows the equipment being used in
User Lab 3 to be operated from the control room via EPICS.  Response and delay time can then be
programmed allowing a more accurate and repeatable measurement.

The new Beamviewer control boxes are back from fabrication.  Assembly of the beamviewers and control
boxes will begin soon.

Studies were done cold cathode vacuum controllers for both the upgrade and the SNS project. These gauges will
be the vacuum interlock for the input couplers. The absolute pressure and time response to trips were measured,
these can be seen at:  http://laser.jlab.org/sns.

The Laser web site upgrades continue.  A new format is being incorporated that will included photos from the
Photo Gallery on the home page so people will be able to see some of what we are doing as soon as they log on.
The task list continues to become more functional as well as minor improvements are made to make it more user

The design phase of the Digital Isolation Module and the TTL Fan-Out Module is complete.  The TTL Fan-Out
Module drawing package is also complete and these boards are ready to be sent out for production.  Minor
modifications have been made to the General Purpose Carrier Board design and the Programmable Time and
Logic Board design.  EECAD is modifying the drawings to include the changes for the next production cycle.

Welcome to our newest I & C Team member, Jim Coleman who comes to us from the safety group.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
•   Consulted with Magnet Enterprises International on the details of the contract while the material is being prepared
    for a shipment next week
•   Bulk steel, copper and pole tip sheet steel were received by Magnet Enterprises International.
•   We plan to inspect the first coil potting Mid November.
Injector Dipoles (DU/DV)
•   The drawings went through final review. The details are fine.
•   Some small portions of the final assemblies need modification.   DULY Research is revising them.
Arc Dipoles (GY, GX, GQ)
•   All steel necessary for the dipoles was received.
•   Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ)
    with the comments and corrections.
•   Effort concentrated on bringing the faces and the gap in conformance with the magnetic model.
•   Dave Douglas is looking at the results of the magnetic model in anticipation of approving them.
•   The drawings and manufacturing specifications for the 180° Dipole (DY) were signed and the package was
    sent to Procurement.
•   The Procurement Department sent the package for the 180° Dipole (DY) out to vendors in the Department’s
    first totally electronic solicitation.  The drawings and specifications are in a WEB site available to all vendors for
    download.  This allows a better search for best value because of the expanded number of vendors allowed to
    participate.  We eliminated sending the cumbersome 3/4 inch stack of drawings and 3/4 inch pack of
    specifications and schedules.
•  We notified a number of manufacturers throughout the country about the coming procurement and made
    preliminary drawings and specifications available to them via the Web.
Magnet Measurement
•   The drawings for the measurement stand for the two quadrupoles were signed and released for manufacture
    and are in a procurement cycle.
•   The design of the dipole measurement stand continued with the all important probe drawing nearing
    completion.  A final resolution of the measurement of the GX dipole’s UV branch held up final approval.   We
    clarified the results of the magnetic model with AES to optimize design of the probe for the GX dipole’s UV
•   We started work on the probe that will measure the sextupole and the octupoles
3 inch quad (QX)
•   The assembly was awarded to New England Technicoil, the core vendor.
•   Al Guerra inspected the machining of the first core at New England Technicoil.  He resolved an interpretation
    of the drawings on the exact nature of the pole tip shape.
•   The first batch of rough-machined cores had gone through annealing.
•   New England Technicoil moved up their schedule for shipment of the first batch from December to November.
•   64 coils are ready to ship from Magnet Enterprises International to the assembly vendor.
Trim Quad (QT)
•   This magnet remains in a procurement cycle.
•   The bid for this magnet went back out for a quick bid cycle to make available Jefferson Lab supplied steel.
    Some vendors were having difficulty getting that grade.
•   Those proposals from the quick re-bid cycle for this magnet are being evaluated
Sextupole (SF)
•   Work on the magnetic model was interrupted because of cooling problems in the septum magnets of the
    CEBAF Accelerator.
•   We have decided to go to DULY Research to get the 3D magnetic model of this magnet advanced.  Robin
    Wines has other duties that prevent her from making much headway.
•   We will purchase one of the extra slabs from Bethlehem Steel left over from the dipole steel order for the solid
    cores of these sextupoles.
Octupole (OT)
•   DULY Research is created a proposal to respond to a Task Order on the Engineering Services Contract to
    convert the magnetic model into a full blown magnetic design.  That proposal was awarded.
•   DULY then completed their preliminary mechanical model of the octupole with an excellent analysis of the
    required shape of the pole tip and field clamp to achieve David Douglas’ specifications.  They have now been
    diverted to Light Box design and the Octupole is put on hold because it is less critical.
Corrector Dipoles (DB, DJ)
•   Milhous Co. has made approximately 20% of the coils.  They shipped the first batch.  They are late but the
    coils look very good.  They are under notification to complete the contract by Thanksgiving.
Beam Line and Vacuum
•   The features of the chamber in the for the injection-reinjection beam line chambers were settled.
•   The modifications of the injector line stands and girder for dipoles were finalized.
•   The girders and pedestals necessary for the return leg and the hardware for the entire machine remain in a
    procurement cycle.
•   Dave Waldman generated the preliminary task order for design of the Arc Chambers utilizing the Engineering
    Services contract with Advanced Energy Systems and they have given us feed back.
•   Design work on girders and stands continued at a high level on the quadrupole telescope in front of the first arc
    and the region between the cryomodules.   A new stand concept using three posts out of one base for these
    congested areas is being incorporated in all layouts.
• Design of the LCW system started up again with presentation of the layout to the Monday Design Meeting

WBS 10 (Wiggler):
After some detective work, Tommy Hiatt and Michael Necaise found that the motor controller for the magnet
measurement stand was causing a 2.5mV offset on the probe.  This leads to the 5 G offset in the probe noted in
last weeks report.  The end corrector strength used last week was found to be the correct one.  Hiatt and
Necaise also found that a strange variation in the sinusoidal component in the vertical field was due to a 0.4
degree twist in the rail supporting the probe.  Since the horizontal field is so uniform however, it does not lead to
a large alteration of the horizontal trajectory.

By rearranging the magnet test setup and using a shorter cable, M. Necaise was able to greatly reduce the noise
leading to DC offsets in the magnet measurement setup. Production runs have started for wiggler field
measurements.  We were able to reduce the level of noise in the system substantially by adding a grounded
sheath to the Hall probe cable.  So far, the wiggler quality looks excellent with the wiggler exceeding its
specifications nicely.

The vacuum chambers are complete and leak checked with the exception of the two flanges which mate to the
dispersion section.  These will be welded when fitup is accomplished.  The aperture of the chamber was
checked and it was found adequate for operation out to 25 microns if a Rayleigh range of 340 cm is used and
the cavity waist is in the center to within 30 cm.

WBS 11 (Optics):
Procurements continue to be made on optical cavity components.  These, and other components have
transitioned to detailing.  We are working on other components such as the insertable mirror assembly and the
active  stabilization systems.  Copies of the presentation made at the internal review in September were sent to
C. Brau and J. Albertine for review and comments.  Particular attention is being devoted to the sizing of the
vertical translation rails in the optical cavity vacuum chambers.  After the internal review, we thought they
needed to be resized, but this would have considerably changed component placement.  With some changes in
rail separation and component stack up, the design team came up with a solution that drops the rail diameters to
the original size(due to reduced torque).   They have also slightly redesigned the yaw mount to make it stiffer.
We learned that the lab has had some problems with using Dicronite (tungsten disulfide) as a high
vacuum-compatible low-friction coating, and so we're looking at other options.

The upgrade near-Brewster windows were inspected, cleaned, and one was installed in user lab 1. New IR
cameras made by Raytheon are being installed in several locations in the optical transport, as well as on the
larger aperture pick off provided by the upgraded near-Brewster window in lab 1.  These cameras will give us
information on the beam position.

The protected-silver coated moly mirror in the present collimator that absorbs the most power was found
discolored.  This is a failure mode associated with running at high average power for extended periods of time,
as was done for the JTO run.  In an attempt to see whether we can mitigate this, we've switched to a
protected-silver on silicon mirror.  Meanwhile, the two damaged mirrors are being sent out to be polished and

Michelle Shinn attended the Boulder Damage Symposium, the premier conference devoted to optical damage in
laser components.  Two highlights that positively impact the upgrades to higher power are:
(1) Magnetorheological finishing greatly reduces subsurface damage introduced during fabrication, with
concomitant increases in damage threshold and large reductions in damage probability near threshold.
(2) New measurements on absorption losses in commercial grade sapphire are lower by an order of magnitude
than originally modeled.

We established operation of the FEL autocorrelator and got good measurements of the optical pulse length for
comparison with predictions from the spectral width.  We intend to continue these experiments parasitically to
get the measurements as a function of FEL detuning.

We provided support for our FEL experiments, specifically, measuring the beam quality as a function of power
using diagnostics installed in the vault, installing a 3 micron high reflector as the outcoupler in order to enable
second harmonic lasing, measuring the polarization of the FEL 2nd harmonic laser emission, and measuring the
pulsewidth of the laser when operating at the fundamental (high efficiency lasing).  We also tried operating the
laser at high intracavity power in an attempt to create laser-induced damage to the coatings.  We may have
done it; we saw a spike in the pressure near both mirrors, and an increase in outcoupling.

Using the FEL we continued studies of mirror thermal distortion  as performed last summer.  The results are
under analysis but it appears the mode movement which leads to an increase of M^2 is a function of power and
actually initially decreases as the power is raised. We spent one day on second harmonic lasing of the FEL as a
continuation of work from last January in hopes of providing a better explanation of the generation mechanism.
We uncovered further mysteries: the donut mode was found to have linear polarization instead of the expected
radial.  We then spent two days on studies aimed toward high FEL efficiency.  We got results which clearly
showed turbulent behavior but we could not duplicate the very high efficiencies seen by the Japanese.  Finally
we attempted to find the damage limits of the FEL mirrors and increased the cavity power in a very high Q
setup.  The FEL continued lasing under all conditions tried and the mirrors are presently being removed to see if
minor damage occurred. None was seen.

We welcome Dr. Shukui Zhang to the FEL Dept.  He will be working in the Optics Group on new diagnostics
for the FEL.

We began the Fall User run by re-establishing beam from the gun after a heat clean cycle.  The high voltage
processing went poorly with frequent arcs.  It was later discovered that the conditioning resistor had not been in
the circuit, probably due to a short, and we believe this has led to minor damage on the cathode which is
presently limiting our operating voltage to 300 kV or below.  We have established operation at this voltage and it
is stable. We then performed a set of nanotube production experiments producing significant quantities
(significant fractions of a gram) of nanotubes for analysis.

The first attempts by members of the UVa Chemistry Dept. to look at vibrational excitation of molecules in a
molecular beam were executed this week.  Using bolometric detection of molecules of pyrone in a supersonic
nozzle beam, a strong signal was observed. We are currently focusing on signal to noise issues.  A Michelson
arrangement will later be used to generate 2 beams with a variable time delay for intramolecular dynamics

An additional effort at near field microscopy at 3 microns was attempted but was unsuccessful.
Work on Thompson X-ray studies continued as a background study.